OpenStudy (anonymous):

find the 9th term of the sequence described by : A(n)=7+(n-1)(3) A.42 B.9 C.31 D.27

OpenStudy (anonymous):

@inowalst @QemarGray6 @Love_Ranaa @Conqueror @Abhisar

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_{n}=7+3(n-1) }\) is the formula for any \(\large\color{black}{ \displaystyle n }\)th term in the sequence. For instance, if I wanted to find the 5th term, I would do the following. \(\large\color{black}{ \displaystyle a_{5}=7+3(5-1)=7+3(2)=7+6=13 }\) So, \(\large\color{black}{ \displaystyle a_{5}=13 }\)

OpenStudy (anonymous):

so would it be B ?? @SolomonZelman

OpenStudy (solomonzelman):

My formula is \(\large\color{black}{ \displaystyle a_{n}=7+3(n-1) }\). just as I plugged in 5 for n to find the 5th term, so would I plug in 9 for n to find the 9th term. it is not B.

OpenStudy (anonymous):

D

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Alright, can you plug in 9 for n, into: \(\large\color{black}{ \displaystyle A_{n}= 7~+~3(n-1) }\) ?? \(\Large\color{black}{ \displaystyle A_{\color{red}{9}}= 7~+~3(\color{red}{9}-1) }\)

OpenStudy (solomonzelman):

then, all you have to do is to simplify the right hand side of the question.

OpenStudy (solomonzelman):

of the *equation*

OpenStudy (anonymous):

i got A .:)

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

1) what is 9-1? 2) multiply that times 3, what do you get after this? 3) Then add 7 to the number that you got from 2 previous steps.

OpenStudy (anonymous):

31!!!!!!!!!